OpenStudy (anonymous):

Solve the equation. Check your solution 8v - 3v - 3 = 22

OpenStudy (aroub):

5v-3=22 5v=19 v=19/5

OpenStudy (anonymous):

How do i do this though, like what are the steps??

OpenStudy (aroub):

ok its really easy you have 8v and -3v which is 8v-3v=5v and that will make 5v-3=22 and then u have to find v right so u add 3 on the both sides to get rid of it so it becomes 5v=19 then u add 1/5 on both sides to get rid if the 5 so then it becomes v=19/5 hope u understand :)

OpenStudy (anonymous):

i wish there was voie chats on here im so confused like how did you get the 5v

OpenStudy (anonymous):

nevermind i get it, thanks!!

OpenStudy (aroub):

ya i agree with u about the voice chat and its ok i will tell u look whats 8-3 its 5 and they have a common variable which is v so it becomes 5v :)